### v1

# LanguageBind: Extending Video-Language Pretraining to N-modality by Language-based Semantic Alignment

Abstract
The video-language (VL) pretraining has achieved remarkable improvement in multiple downstream tasks. However, the current VL pretraining framework is hard to extend to multiple modalities (N modalities, $N\geq 3$) beyond vision and language. We thus propose LanguageBind, taking the language as the bind across different modalities because the language modality is well-explored and contains rich semantics. Specifically, we freeze the language encoder acquired by VL pretraining and then train encoders for other modalities with contrastive learning. As a result, all modalities are mapped to a shared feature space, implementing multi-modal semantic alignment. While LanguageBind ensures that we can extend VL modalities to N modalities, we also need a high-quality dataset with alignment data pairs centered on language. We thus propose VIDAL-10M with 10 Million data with Video, Infrared, Depth, Audio and their corresponding Language. In our VIDAL-10M, all videos are from short video platforms with complete semantics rather than truncated segments from long videos, and all the video, depth, infrared, and audio modalities are aligned to their textual descriptions. After pretraining on VIDAL-10M, we outperform ImageBind by 1.2% R@1 on the MSR-VTT dataset with only 15% of the parameters in the zero-shot video-text retrieval, validating the high quality of our dataset. Beyond this, our LanguageBind has greatly improved in the zero-shot video, audio, depth, and infrared understanding tasks. For instance, on the LLVIP and NYU-D datasets, LanguageBind outperforms ImageBind-huge with 23.8% and 11.1% top-1 accuracy. Code address: https://github.com/PKU-YuanGroup/LanguageBind
1 Introduction
With the development of the Internet and smartphones, there has been a proliferation of video websites and apps (e.g., Youtube and TikTok), leading to a substantial increase number of videos (Xue et al., 2022). Consequently, a set of video tasks have emerged, such as video search (Smith & Chang, 1997), video recommendation (Deldjoo et al., 2016), and video editing (Casares et al., 2002; Bonneel et al., 2014). To solve video understanding tasks, video-language pretraining has been employed by training foundation models by combining computer vision (He et al., 2016; Dosovitskiy et al., 2020) and natural language processing (Vaswani et al., 2017). These models can capture video semantics and solve downstream tasks (Karpathy et al., 2014; Mithun et al., 2018).
However, current VL pretraining frameworks are often limited to vision and language modalities. The ImageBind (Girdhar et al., 2023) introduces an indirect alignment method for multi-modal pretraining. It aligns other modalities to images, facilitating a comprehensive understanding of various modalities such as infrared (Jia et al., 2021), depth (Kim et al., 2022), audio (Piczak, 2015), and IMU (Grauman et al., 2022). In practical tasks such as zero-shot retrieval and classification as shown in Figure 1, the alignment with language modality is predominantly required for various modalities. While the indirect alignment of ImageBind may result in performance degradation, the LanguageBind method does not need images as intermediaries and facilitates straightforward expansion to additional modalities in downstream tasks.
In this paper, we propose the LanguageBind, a language-based multi-modal pretraining framework that can extend video-language pretraining to multiple (N) modalities. As the language modality contains rich semantic information and is well-explored (Kenton & Toutanova, 2019; Dai et al., 2019), we take it as the bind across different modalities. This process maps all modalities to a unified embedding space, enabling effective semantic alignment. To improve training efficiency, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021) for fine-tuning, achieving impressive training results with minimal training iterations.
Figure 1: ImageBind vs. LanguageBind. The ImageBind method relies on images as intermediaries, while the LanguageBind method dispenses with this requirement. LanguageBind directly aligns all modalities to the language space, thereby enhancing its applicability to downstream tasks. “X” represents all modalities except language, and “c” represents category.
To further improve the modal integrity in pretraining and validate our LanguageBind, we introduce a dataset with five modalities, the VIDAL-10M, which includes VL, IL (infrared-language), DL (depth-language), and AL (audio-language) data pairs. The videos of previous datasets are always truncated segments from long videos (Miech et al., 2019; Xue et al., 2022), resulting in fragmented semantics. To avoid this problem, we construct our video-text pairs from short videos with complete stories. To ensure the quality of the central language modality, we perform multi-view text generation and enhancement on VIDAL-10M.
The proposed LanguageBind ensures that we can extend vision-language to multiple (N) modalities, and our dataset VIDAL-10M benefits more downstream tasks beyond VL tasks, including video retrieval (Luo et al., 2022), depth classification (Cao et al., 2017), infrared classification (Baffa & Lattari, 2018) and audio classification (Palanisamy et al., 2020). In text-to-video zero-shot retrieval on the MSR-VTT dataset, LanguageBind outperforms ImageBind with a 1.2% higher R@1 score, while using less than 15% of the parameters. On the MSVD dataset, LanguageBind achieves a notable R@1 score of 42.7%, surpassing the previous state-of-the-art model, CLIP4Clip (Luo et al., 2022), with an improvement of 4.2%. For zero-shot classification on depth and infrared data, LanguageBind achieves a substantial performance advantage over ImageBind. LanguageBind attains top-1 accuracy of 87.2% and 65.1% on LLVIP and NYU-D, respectively, outperforming ImageBind by 23.8% and 11.1%. For zero-shot audio classification tasks, LanguageBind outperforms ImageBind with a 2.7% higher top-1 accuracy on the ESC50 dataset.
We summarize our primary contributions as follows:
  • We propose LanguageBind, the langauge-based multi-modal pretraining approach. During the pretraining process, all modalities gradually align with the language modality through contrastive learning, and these modalities are unified within a shared embedding space.
  • We introduce VIDAL-10M, a large-scale five-modal video dataset, containing 10 million data pairs with aligned VL, IL, DL, and AL. To the best of our knowledge, VIDAL-10M is the first large-scale video dataset with depth and infrared modalities.
  • Extensive experiments validate the effectiveness of our dataset and approach, achieving remarkable performance in video and other modality understanding tasks.
2 Related work
Multi-modal Pretraining
Multi-modal pretraining begins with pretraining in vision and language. CLIP (Radford et al., 2021) pioneered the alignment of images and texts on a large-scale dataset comprising 400 million samples, effectively establishing a bridge between the image and text domains. This alignment benefits a variety of downstream tasks, including zero-shot classification and image-text retrieval. CLIP can also be used as a foundation for alignment in other modalities. For instance, CLIP4Clip (Luo et al., 2022) aligns video with text, CLAP (Wu* et al., 2023) aligns audio with text, and PointCLIP (Zhang et al., 2022) aligns point clouds with text. Recent efforts have undertaken a comprehensive exploration of multi-modal alignment through pretraining. Augmenting the alignment process with additional modalities can enhance the model’s robustness while maintaining its performance, as observed in VALOR (Chen et al., 2023a) and VAST (Chen et al., 2023b). However, as the number of modalities increases, the training paradigm required to align them effectively undergoes significant changes. Meta-transformer (Zhang et al., 2023) accommodates 12 modalities and utilizes distinct tokenizers to harmonize the embedding space across modalities. ImageBind (Girdhar et al., 2023) expands multi-modal alignment pretraining to encompass six modalities but may not perform as well in language-related tasks due to indirect alignment. In our work, we propose LanguageBind, a direct alignment mechanism designed to align alternative modalities directly with the language modality, which has the highest information density. This direct alignment mechanism yields discernible improvements in downstream task performance.
Table 1: Comparision of existing multi-modal datasets. VIDAL-10M is the currently first accessible multi-modal dataset including aligned VL, IL, DL, and AL data pairs.
Datasets Samples Modality Year
HMDB-51 7K V 2011
UCF-101 13K V 2012
ActivityNet-200 20K VT 2015
WebVid-10M 10.7M VT 2021
HD-VILA-100M 100M VT 2022
HowTo-100M 136M VT 2019
LLVIP 15k VI 2021
FLIR 10k VI 2015
FLIRV2 12k VI 2015
NYU-D 1.4k VD 2012
YouTube-8M 6.1M VAT 2016
AVA 58K VAT 2017
VIDAL-10M (Ours) 10M VIDAL 2023
Multi-modal Datasets
Multi-modal datasets serve as the foundation for multi-modal pretraining. Initially, these datasets only consisted of videos and their corresponding categories, as shown in Table 1. HMDB-51 (Kuehne et al., 2011) and UCF-101 (Soomro et al., 2012) are examples of such datasets, which contain truncated segments from long videos with manual annotation. However, creating these datasets required significant human effort, which limited their scalability and diversity. To address this issue, researchers turned their attention to the abundance of video-text resources available on the internet. Inspired by the success of image-text datasets (Sharma et al., 2018; Changpinyo et al., 2021), they used script-based programming (Schuldt et al., 2004; Kong et al., 2019; Sigurdsson et al., 2018) to extract millions of video-text data pairs. However, acquiring data from modalities like infrared (Teledyne FLIR, 2015a; b) and depth (Silberman et al., 2012), which require specialized equipment and manual annotation, has been challenging. This has severely limited the scale of the data and its alignment with other modalities. Although existing work like ImageBind (Girdhar et al., 2023) has attempted to bind various image-paired datasets and achieve indirect semantic alignment between different modalities, this approach still faces issues of incomplete and indirect data alignment. Thus, there is an urgent need for multi-modal datasets with direct semantic aligned data pairs, especially for modalities with five or more types.
Figure 2: LanguageBind overview. The language encoder parameters are frozen, while the multi-modal encoder parameters can be adjusted using the LoRA technique. By employing contrastive learning between language and other modalities, LanguageBind successfully achieved multimodal joint learning, thereby fostering semantic alignment across different modalities.
3 Method
In this section, we present LanguageBind, a multi-modal pretraining approach designed to align the semantics of different modalities and enhance cross-modal retrieval and zero-shot classification. As shown in Figure 2, LanguageBind consists of three parts: (a) multi-modal encoders, (b) language encoder, and (c) multi-modal joint learning.
3.1 Multi-modal Encoders
For the video encoder, we adopted the training framework of CLIP4Clip. The video encoder is initialized from ViT-B/32 and more details are in Appendix B. For other modalities besides language, we employ the 24-layer, 1024-dimensional vision transformer with a patch size of 14. The encoders are initialized from OpenCLIP-large (Ilharco et al., 2021). Depth and infrared are treated as RGB images, which are replicated 3 times in the channel dimension to align with RGB images. Following ImageBind, audio data is transformed into spectrograms with a duration of 10 seconds (128 mel-bins) and we repeat and pad the spectrograms. For example, a 4-second spectrogram would be repeated twice and then padded with zero for an additional 2 seconds. Similarly, we also replicate it 3 times in the channel dimension. If the duration exceeds 10 seconds, we randomly sample three 10-second audio segments, each from the front 1/3, middle 1/3, and back 1/3 of the original audio, and finally stack them together.
Patch masking
To address the inefficiency in processing all tokens within the encoder, we divide the image into patches and take a small portion of patches by encoder mask $\mathbb{M}_{e}$, following MAE (He et al., 2022). Given a modality ${\displaystyle{\bm{m}}}\in\mathbb{R}^{H\times W\times C}$, where $(H,W)$ represents the resolution of the original data, with $C$ denoting the number of channels. We first transform it into patches using a patch embedding layer with non-overlapping filters. This operation produces patches denoted as ${\displaystyle{\bm{m}}}^{\prime}\in\mathbb{R}^{N\times C}$ and $N=\frac{H\times W}{S^{2}}$ represents the resulting sequence length, where $S$ represents the size of each patch. Subsequently, positional embedding is applied to the visible tokens, which are divided by encoder mask. The combined sequence $\bm{x}$ is defined as:
${\displaystyle{\bm{x}}}=\{{\displaystyle{\bm{m}}}^{\prime}_{i}+{\displaystyle{\bm{P}}}_{i}\}_{i\in\mathbb{M}_{e}}$ (1)
where ${\displaystyle{\bm{P}}}$ is a sequence of learnable position tokens, and $i$ represents the position index at patches.
LoRA fine-tuning
We employ the LoRA technique  (Hu et al., 2021) to accelerate fine-tuning. For a modality-agnostic encoder with a weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we maintain the weight matrix $W_{0}$ frozen while learning a new weight matrix $BA$. For instance, in the case of the modality-agnostic encoder $h(\cdot)$ and $\bm{x}$, the forward process can be represented as follows:
$h({\displaystyle{\bm{x}}})=W_{0}{\displaystyle{\bm{x}}}+BA\bm{x}$ (2)
where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, with $r$ being the minimum of $d$ and $k$. It is important to highlight that both $W_{0}$ and $BA$ possess the same input and output dimensions, facilitating their summation to produce the final output.
Modality extending
To extend the LanguageBind method to multiple (N) modalities, the first step involves processing the data into a sequence of tokens. Subsequently, the parameters are initialized from OpenCLIP. The encoder for different modalities is then trained through token masking and LoRA fine-tuning while keeping the language encoder frozen. Finally, this modality is aligned with the language feature space.
3.2 Language Encoder and Multi-modal Joint Learning
For the language encoder, we utilize a 12-layer transformer model with 768-dimensional and initialize it from OpenCLIP. For a given text, we initially employ a BPE tokenizer to segment words into relatively common subwords. Each subword corresponds to a unique token, and these tokens are embedded within a word embedding layer. Ultimately, the tokens are encoded by the language encoder to obtain a text logit $\bm{y}\in\mathbb{R}^{L\times C}$, where $L$ represents the length of the sequence. To ensure alignment across different modalities, we implement contrastive learning principles  (Radford et al., 2021). The objective of this approach is to increase the similarity of paired data, bringing them closer to the same semantic space, while minimizing the similarity of unpaired data. We utilize contrastive learning to bind individual modalities to language.
$L_{M2T}=-\frac{1}{K}\sum_{i=1}^{K}\log\frac{\exp(x_{i}^{\top}y_{i}/\tau)}{\sum_{j=1}^{K}\exp(x_{i}^{\top}y_{j}/\tau)},L_{T2M}=-\frac{1}{K}\sum_{i=1}^{K}\log\frac{\exp(y_{i}^{\top}x_{i}/\tau)}{\sum_{j=1}^{K}\exp(y_{i}^{\top}x_{j}/\tau)}$ (3)
where $x_{i}$ is the $i$-th modality data and $y_{j}$ is the $j$-th text and their features are normalized. $K$ and $\tau$ are batch size and the temperature. The direct alignment of each modality $\mathbf{M}$ with language $\mathbf{T}$ enables us to significantly enhance zero-shot classification and retrieval tasks.
4 The VIDAL-10M dataset
In this section, we will describe how to construct our VIDAL-10M dataset, including 3 million pairs of video-language data, 3 million pairs of infrared-language data, 3 million pairs of depth-language data, and 1 million pairs of audio-language data. As shown in Figure 3, the collection process consists of three main steps: visual search term database construction (Section 4.1), video and audio collection and filtering (Section 4.2), and modality generation and enhancement (Section 4.3).
Figure 3: VIDAL-10M construction. (a) Firstly, a search term database is generated by leveraging visually related datasets. (b) Subsequently, relevant videos and audios are collected from the internet and undergo a series of filtering processes. (c) Lastly, we perform infrared and depth modality generation, as well as multi-view text generation and enhancement.
4.1 Visual search term database construction
To build a video dataset with rich visual concepts and diversity, we design a unique search term acquisition strategy. We leverage text data including labels and captions from various visual task datasets (YouTube-8M (Abu-El-Haija et al., 2016), MSR-VTT (Xu et al., 2016), COCO (Lin et al., 2014), AVA (Gu et al., 2018), HMDB-51 (Kuehne et al., 2011), ImageNet (Deng et al., 2009)) to create a large-scale search term database with diversity and broad applicability. Then we filter these search terms based on their frequency and employ the NLTK toolkit for part-of-speech tagging, followed by tallying the occurrences of keywords (nouns and verbs). A balanced subset of 100,000 search items corresponding to these keywords is then extracted as the final search team database.
4.2 Video and Audio collection and filtering
During the data collection process, we utilize the aforementioned search terms to retrieve video-text pairs and audio-text pairs from relevant platforms, e.g. YouTube Shorts, and Freesound. Regarding video collection, in order to obtain short videos with high-quality textual descriptions, we implemented a filtering mechanism for the title and hashtags. Video samples with titles containing less than 2 words and without video hashtag labels are excluded from our dataset. Moreover, we removed irrelevant words and hashtags, such as ”youtube”, ”fyp”, ”shorts”, etc. Furthermore, to ensure a complete, consistent, and precise depiction of the event within a single full video, we decide to impose a duration limit of 20 seconds. Shorter videos tend to exhibit better scene coherence and event integrity and are more closely aligned with corresponding hashtags and title descriptions. Ultimately, we obtain a short video dataset that encompasses more specific rather than abstract content. Concerning audio collection, we rank the audio list on the Freesound platform based on its similarity to the search terms. Additionally, we conduct filtering operations similar to those for videos, taking into account factors such as audio ratings, download counts, user comments, tags, and duration. This comprehensive approach allows us to curate and refine the audio and video content more effectively.
Figure 4: Multi-view text generation and enhancement pipline. We employ the OFA model to generate keyframe captions and input video, title and hashtags into the mPLUG-owl model to obtain video captions. The video captions are further refined using ChatGPT, resulting in the ChatGPT-mPLUG caption. The final multi-view textual description comprises these components.
4.3 Modality generation and enhancement
Multi-view text generation and enhancement
The language modality of VIDAL-10M consists of multi-view texts, including title, hashtags, keyframe captions, video captions, and enhanced captions. The detailed text generation and enhancement pipeline is illustrated in Figure 4. Hashtags in VIDAL-10M are specifically designed to highlight the main subjects and actions depicted in the video. These hashtags serve as key indicators, emphasizing the focal points and dynamic elements of the video. However, hashtags alone may not fully capture the spatial information conveyed by the video frames. To address this limitation, we leverage the image captioning model OFA (Wang et al., 2022) to generate supplementary keyframe captions that enrich the spatial information at the keyframe level. These captions also contain local temporal information related to the video content, which is beneficial for visual-text pretraining. Besides spatial information, temporal information concealed within the video is equally significant, providing crucial insights into the progression and sequencing of events within the video. To further supplement the overall thematic and temporal information of the video, we employ the mPLUG-owl model (Ye et al., 2023) to generate video captions based on the combination of video, title, and hashtags. By leveraging the title and hashtags as accurate video labels, we guide the mPLUG-owl model to generate captions that align with the video theme, reducing potential model bias to a certain extent. Furthermore, to extract valuable information from the generated video captions, we utilize the ChatGPT model to refine and enhance the textual description, thereby greatly improving the quality of the text. By incorporating the above text components, multi-view textual descriptions provide a comprehensive and detailed representation of the video content.
Infrared and depth modality generation
In the field of depth and infrared, creating modal datasets typically requires specialized equipment and human effort, resulting in limited data. Despite the success of large-scale pretraining models (Radford et al., 2021; Wu* et al., 2023; Luo et al., 2022; Chen et al., 2023b) in NLP and CV, there remains a lack of large-scale data in this field. To address this challenge, we propose using advanced generative models specifically to construct a large-scale dataset of depth and infrared. The sRGB-TIR model (Lee et al., 2023) is used for infrared modality generation and the GLPN model (Kim et al., 2022) for depth modality generation, generating depth and infrared from keyframes in our videos. While some limitations may exist, our collection of millions of video frames and corresponding texts with highly diverse semantics can significantly reduce the presence of model biases.
5 Experiments and Results
In this section, we evaluate the effectiveness of LanguageBind in various downstream tasks through different experiments. Firstly, LanguageBind’s capability to align video and text is assessed using zero-shot video-text retrieval. Additionally, we use LanguageBind to enhance the performance of downstream tasks that involve depth, infrared images, and audios. Finally, we conduct ablation experiments to analyze the impact of different parameter configurations and text descriptions on LanguageBind’s performance.
5.1 Zero-shot classification in multiple modalities
We compare our model with the recent state-of-the-art multi-modal pretraining models, OpenCLIP (Ilharco et al., 2021) and ImageBind (Girdhar et al., 2023) on multi-modal understanding ability tasks in Table 5.1. For infrared-language zero-shot classification, LanguageBind exhibits a noteworthy 23.8% performance advantage over ImageBind on the LLVIP and outperforms OpenCLIP on all three datasets(LLVIP, FLIR V1, and V2). For depth images, our zero-shot results on NYU-D surpass ImageBind by a substantial margin of 11.1% and outperform OpenCLIP by 19.7%. For audio, we outperform ImageBind by 2.7% on the ESC-50 dataset.
[…]
tableInfrared-Language, Depth-Language, and Audio-Language zero-shot classification. We report the top-1 classification accuracy for all datasets. Method Parameter Infrared Depth Audio LLVIP FLIR V1 FLIR V2 NYU-D ESC-50 Random - 50.0 25.0 8.3 10.0 2.75 ImageBind 633M 63.4 - - 54.0 66.9 OpenCLIP 305M 82.2 81.2 42.6 45.4 - LanguageBind 305M 87.2 82.9 48.0 65.1 69.6
5.2 Zero-shot retrieval in Video-Language
We conduct video-text retrieval experiments on two classical datasets, MSR-VTT (Xu et al., 2016) and MSVD (Chen & Dolan, 2011). To ensure a fair comparison, we utilized the CLIP4Clip training framework and provided additional details in Appendix B. Our method outperformed existing techniques obviously, as indicated in Table 2. For the text-to-video task, we achieved a higher R@1 of 1.2% than ImageBind with fewer than 15% parameters on MSR-VTT. Moreover, we surpassed CLIP4Clip (Luo et al., 2022) by 5.0% and CLIP-straight (Portillo-Quintero et al., 2021) by 5.8% on MSR-VTT, and by 4.2% and 5.7% on MSVD, respectively. Furthermore, we achieved state-of-the-art performance in video-to-text retrieval tasks, significantly outperforming CLIP-straight.
Table 2: Zero-shot Video-Text Retrieval Performance on MSR-VTT and MSVD datasets. ”R@n” represents recalls at different points, and ”MR” stands for Median Rank, following the methodology from  (Zhu & Yang, 2020; Luo et al., 2020). We focus on reporting the parameters of the vision encoder. Our experiments are based on 3 million video-text pairs of VIDAL-10M.
Dataset Method Parameter Pt-data R@1normal-$\uparrow$ R@5normal-$\uparrow$ R@10normal-$\uparrow$ MRnormal-$\downarrow$
MSR-VTT SupportSet 78M IN1K, IG65M, HT100M 8.7 23.0 31.1 31.0
HD-VILA 60M HD-VILA-100M 14.6 34.4 44.1 15.0
CLIP-straight 86M WIT400M 31.2 53.7 64.0 4.0
CLIP4Clip 86M WIT400M, HT100M-380k 32.0 57.0 66.9 4.0
(Text-to-Video) HiTeA 37M IN21K, CC15M, WV2.5M 34.4 60.0 69.9 -
ImageBind 633M LAION2B 36.1 - - -
Ours 86M WIT400M, VIDAL-10M 37.3 62.9 71.5 3.0
MSVD SupportSet 78M IN1K, IG65M, HT100M 28.4 60.0 72.9 4.0
(Text-to-Video) Frozen 86M CC3M,WV2.5M 33.7 64.7 76.3 3.0
CLIP-straight 86M WIT400M 37.0 64.1 73.8 3.0
CLIP4Clip 86M WIT400M, HT100M-380k 38.5 66.9 76.8 2.0
Ours 86M WIT400M, VIDAL-10M 42.7 71.2 80.5 2.0
MSR-VTT CLIP-straight 86M WIT400M 27.2 51.7 62.6 5.0
(Video-to-Text) Ours 86M WIT400M, VIDAL-10M 37.1 62.6 71.5 3.0
MSVD CLIP-straight 86M WIT400M 59.9 85.2 90.7 1.0
(Video-to-Text) Ours 86M WIT400M, VIDAL-10M 63.9 88.0 93.1 1.0
6 Ablation Study
In this section, we conduct extensive experiments to investigate the impact of several factors. At first, we examine the effects of different enhanced textual inputs on downstream tasks. Furthermore, we assess the impact of data volumes on pretraining. In addition, we explore various training strategies to enhance zero-shot classification. Finally, we conduct a meticulous analysis of model training configurations to ensure robust transferability.
6.1 Impact of different text sources
In Table 6.1, we conduct various experiments to explore how different text sources impact language modality. We verify the effectiveness of LanguageBind, trained with text from multiple sources, across various modalities. While some text sources yield good results, we discover that a single text source may not be universally suitable for all downstream tasks and datasets. In terms of video and depth modalities, the ChatGPT enhanced caption proves to be advantageous. For infrared images, the OFA performs best in the LLVIP dataset, while the raw caption achieves the highest accuracy in FLIR v1 and v2. That’s why our VIDAL-10M provides multi-view textual descriptions, allowing for flexibility in selecting an appropriate text source that caters to diverse task requirements.
[…]
tableImpact of different text sources. We report the results of text-to-video R@1 for zero-shot retrieval and other datasets report top-1 accuracy. MSR-VTT results were tested on a 500K subset of VIDAL-10M. “Raw caption” denotes the title & hashtags. Modality Dataset Raw caption OFA caption mPLUG caption ChatGPT-mPLUG caption Video MSR-VTT 33.5 34.5 35.8 36.4 Infrared LLVIP 83.9 87.2 84.6 84.8 FLIR V1 82.9 80.6 81.4 81.6 FLIR V2 48.0 45.7 46.8 46.6 Depth NYU-D 61.5 62.1 63.9 65.1
6.2 Scaling the size of dataset
We analyze the impact of different data amounts on MSR-VTT and report the R@1 score for zero-shot retrieval as shown in Figure 5. Our findings indicate that an increase in data amount leads to significant improvement in recognition performance. Specifically, the performance of 3M ChatGPT-enhanced text surpasses that of 500k and 100k data by 0.9% and 1.6%, respectively.
Figure 5: Scaling pretraining data size.
Furthermore, the trends observed in both video-to-text retrieval and text-to-video retrieval consistently demonstrate that the interaction between modalities plays a pivotal role in enhancing the learning process. Consequently, with the expansion of data size, the textual descriptions within the VIDAL-10M dataset align more closely with the video content and demonstrate increased scalability.
6.3 Training Loss and Architecture
Following ImageBind, we focus on depth modality, which is visual and spatial modality. We report top-1 accuracy on the NYU-D dataset.
Training epochs
We conduct an experiment in Table 3a to study the effect of training epochs which shows that the LoRA fine-tuning is highly effective. Although 3 epochs of training regimen yield superior accuracy, we chose to optimize for a single epoch, achieving a balance between performance and training cost.
Training batch size
In Table 3b, we evaluate the effect of batch size on representation learning. The experiments have shown that a larger batch size is not necessarily better. In fact, a batch size of 1,024 is the most optimal.
Rank of LoRA
In our investigation, we examined prevalent rank configurations for LoRA, as detailed in Table 3c. We observe that smaller rank values lead to more significant performance improvements, whereas larger rank tends to decrease performance. This trend may be attributed to the potential overfitting of the model.
Temperature for loss
We scrutinized the impact of diverse temperature in Table 3d. We found that the learnable temperature initiated from 0.07 performs best, outperforming the fixed temperature strategy proposed by ImageBind.
Masked ratio
We explore the impact of different mask ratios in Table 3e. The results show that a mask ratio of 0.5 demonstrates the highest performance, requiring only a quarter of the computational resources, aligning with findings in FLIP (Li et al., 2023).
Table 3: Training loss and architecture design decisions and their impact on zero-shot classification. Settings for results in Section 5.1 highlighted in gray.
1 3 5
65.1 65.4 64.5
(a) Epochs
512 1k 2k
63.9 65.1 64.5
(b) Batch size
2 4 8
65.1 64.4 64.7
(c) Rank of LoRA
Learn 0.05 0.1
65.1 63.0 62.7
(d) Temperature
0.3 0.5 0.7
64.8 65.1 62.7
(e) Masked ratio
7 Conclusion
In this work, we propose the LanguageBind, a language-based semantic alignment method for multi-modal pretraining. We employ contrastive learning to establish modality semantic alignment between the language modality and all other modalities. To improve modal integrity, we also construct the first large-scale multi-modal dataset directly aligned to language modality, VIDAL-10M, comprising 10 million aligned VL, IL, DL, and AL pairs. Experimental results on the zero-shot VL, IL, DL, and AL understanding tasks demonstrate the capability of LanguageBind and the effectiveness of VIDAL-10M.
Reproducibility Statement
  1. 1.
    For LanguageBind approach details.
    1. (a)
      We provide a comprehensive overview of the multi-modal encoder, detailing its architecture and functionality in Section 3.1.
    2. (b)
      We outline the language encoder in Section 3.2.
    3. (c)
      We expound on the methodologies employed for multi-modal joint learning in Section 3.2
  2. 2.
    For VIDAL-10M dataset construction details.
    1. (a)
      We describe the procedures employed to construct the search term database in Section 4.1.
    2. (b)
      We provide insights into the strategies used for collecting and filtering video and audio data within VIDAL-10M in Section 4.2.
    3. (c)
      We elaborate on the generation of infrared and depth data, as well as the processes involved in multi-view text generation and enhancement in Section 4.3
    4. (d)
      We promise to release the VIDAL-10M dataset upon publication.
  3. 3.
    For setting details.
    1. (a)
      We describe in detail the training hyperparameters in Appendix B.
    2. (b)
      We describe the setup of the downstream task dataset Appendix C.
Appendix
Appendix A Statistics of VIDAL-10M Dataset
Figure 6: Examples of video-audio-text-depth-infrared pairs in VIDAL-10M, with the text components comprising hashtags, title, keyframe captions, and video caption. Examples are taken from 4 distinct clusters, corresponding to Sports, Pets & Animals, News & Politics, and Education.
Table 4: Examples of textual descriptions from various datasets as search terms.
Dataset Search terms
YouTube-8M How to make a delicious chocolate cake.
Learn to dance salsa in 10 easy steps.
Howto100M How to play chess.
How to make pizza.
ImageNet lesser panda, red panda, panda, bear cat, cat bear, Ailurus fulgens, coon bear
killer whale, killer, grampus, sea wolf, Orcinus orca, giant panda, panda, panda bear
COCO A small boat floating on a body of water with a city skyline in the background.
A man with a red helmet on a small moped on a dirt road.
Others ……
In order to build a video dataset with rich visual concepts and diversity, we develop a unique but simple search term acquisition strategy. This strategy involves obtaining search terms from various visual datasets (as shown in Table 4). Subsequently, we use these search terms to gather videos from the YouTube Shorts platform, which has become a popular source for video data due to its abundance and diverse content. We collect videos in various categories, including sports, animals, nature, etc., resulting in a large and diverse dataset. Examples of video-audio-text-depth-infrared pairs in the VIDAL-10M dataset are shown in Figure 6. Moreover, to ensure data quality, we manually design a list of stop words that are filtered from our datasets. These words include terms such as ”bts”, ”bmw”, and ”nfl”, among others, that are not relevant to our research.
Table 5: Stop words in our datasets.
viral funny love fashion subscribe nature
asmr motivation fitness art satisfying foryou
music india fun bts amazing edit
life roblox vlog minecraft design marvel
explore dubai foryoupage comedy outfit ootd
share indian lol creative relaxing tattoo
random instagram quotes workout sad ideas
views bgmi yummy respect easy usa
ronaldo jawellery memes happy nfl song
mlb reel support nba wow status
gree meme gameplay top blackpink whatsappstatus
follow homedecor history tutorial bodybuilding japan
interiordesign freefire stunt foodie animation recipe
skills tips crazy pov editing aesthetic
style view london reaction story pubg
construction challenge healthy bmw uk free
hairstyle enjoy motivational messi capcut nailart
entertainment fifa attitude europe health geography
gta unboxing adventure whatsapp fail btsarny
god inspiration relatable comment tattoos fy
highlights amazon illustration fortnite ntb avaiation
interior decor travelvlog canada btsarmy tranding
time mtb luxury vlogs picsart reels
photoshoot business photography
Video categories and duration
Furthermore, we analyze the distribution of video categories with varying durations in our datasets, as illustrated in Figure 7. The normal distribution pattern observed in this analysis indicates that our dataset covers a wide range of concepts. Besides, we show the proportions of each category across different duration grades in the VIDAL-10M dataset in Figure 8.
Figure 7: The number of 15 categories with different durations in our VIDAL-10M datasets. A wide range of concepts are covered.
Figure 8: The statistical distribution of categories across the three duration grades in the VIDAL-10M dataset. The colors green, blue, and yellow represent video durations of 1-7, 8-14, and 15-20 s, respectively.
FPS, Aspect ratio and Resolution
The first aspect examined in the dataset is the Frames Per Second (FPS) of the videos. FPS refers to the number of frames or images displayed per second in a video. The aspect ratio of a video represents the proportional relationship between its width and height dimensions. It is a critical factor in determining the visual presentation and viewing experience of the videos. The distribution of FPS and aspect ratios in Figure 9 provides insights into the smoothness and fluidity of the recorded content and sheds light on the various formats and orientations used. Video resolution refers to the number of pixels in each dimension that a video contains. It directly affects the clarity, sharpness, and level of detail in the visual content. Examining the distribution of resolutions (Figure 10) in the dataset provides an understanding of the available video quality and the technological capabilities of the recorded material.
Figure 9: The distribution of FPS (Frames Per Second) and aspect ratio in the videos of the VIDAL-10M dataset.
Figure 10: Height and width distribution of videos in VIDAL-10M dataset.
Appendix B Pretraining details
In this section, we introduce our training configuration.
[…]
tablepretraining setting Config Video Infrared Depth Audio Vision encoder ViT-Base/32 ViT-Huge/14 ViT-Huge/14 ViT-Huge/14 Optimizer BertAdam AdamW AdamW AdamW Optimizer Momentum $\beta_{1},\beta_{2}=0.9,0.98$ Epochs 1 1 1 8 Learning rate 1e-4 1e-4 5e-4 5e-4 Coefficient learning rate 1e-3 Weight decay 0.2 Batch size 512 1024 1024 512 Warmup steps - 200 200 200 Temperature learnable Learning rate schedule cosine decay Slice frame position cut from head frames - - - Max words 32 77 77 77 Max frames 12 - - - Cls header mean-pooling - - - LoRA rank - 2 2 8 LoRA alpha - 16 16 16 LoRA dropout - 0.1 0.1 0.1
Video-Language.
For the video-text retrieval, we verify that the VIDAL-10M dataset is highly aligned. We adopted the training framework of CLIP4Clip, and the model is initialized from ViT-B/32, and the rest of the parameters are the same as the default settings, except for 1 epoch and batch size of 512.
Depth-Language.
The model is initialized from OpenCLIP with a frozen language encoder. For each individual sample, we employ a random selection approach to extract either a depth image from the video sequence. Subsequently, we resize these frames to have a short edge length of 256 units, followed by a central cropping process to attain dimensions of 224×224. Additionally, we tripled the number of channels in both the depth image. The text templates employed for zero-shot classification are sourced from OpenCLIP, with a modification consisting of the substitution of ”photo” with ”depth photo” across all templates. This alteration yields an approximate performance gain of 1%.
Infrared-Language.
Following depth-language, it is worth noting that the text templates corresponding to infrared images retain the ”photo” designation, as no discernible performance improvement is observed from this particular modification.
Audio-Language.
The data are preprocessed as in 3.1. Unlike depth and infrared, spectrograms differ much from the domain of conventional visual images. Therefore, it is not easy to overfit during training, so we increase the training epoch and the rank of LoRA. Additionally, we replace ”the/a photo of” with ”the/a sound of” across all templates for audio zero-shot classification.
Appendix C Downstream datasets
Video-language.
We perform video-text retrieval experiments on 2 datasets. (a) MSR-VTT (Xu et al., 2016) comprises 10K YouTube videos, each paired by 200K captions. In our analysis, we present results based on the 1K-A test subset. (b) MSVD (Chen & Dolan, 2011) consists of about 120K sentences and reports results on test data (670 samples).
Infrared-language.
(a) LLVIP (Jia et al., 2021) constitutes a dataset for pedestrian object detection within the infrared spectrum. Following ImageBind, we extracted all people from the images, designating all other objects as background elements. This process resulted in a dataset comprising 7,622 ‘background’ classes and 7,954 ‘person’ classes, which was subsequently employed for binary classification testing. (b) FLIR v1 (Teledyne FLIR, 2015a) offers comprehensive annotations for both thermal and visible spectrum frames. From the test data, we derived a dataset containing 11,696 images by extracting bounding boxes. This dataset encompasses 4 categories – [’bicycle’, ’car’, ’dog’, ’person’]. (c) FLIR v2 (Teledyne FLIR, 2015b) includes 16,696 images after processing similarly, which were categorized into 12 classes – [’bike’, ’bus’, ’car’, ’hydrant’, ’light’, ’motor’, ’other vehicle’, ’person’, ’sign’, ’skateboard’, ’stroller’, ’truck’].
Depth-language.
We use NYU-v2 Depth-only (NYU-D) (Silberman et al., 2012) to validate by 654 test samples. Through preprocessing, we constrained the depth images to a maximum depth of 10 meters. Following ImageBind, we undertook a category reorganization process, resulting in a total of 10 scene categories.
Appendix D License
Unless explicitly noted otherwise, our released datasets are provided to users under the terms of the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International Public License (”CC BY-NC-SA 4.0”), in conjunction with the additional terms outlined herein. The CC BY-NC-SA 4.0 license can be accessed at https://creativecommons.org/licenses/by-nc-sa/4.0/legalcode. By downloading or utilizing our datasets from our website or other sources, you agree to adhere to the terms of CC BY-NC-SA 4.0, as well as the terms outlined in our dataset Terms. In the event of any conflict between the terms of CC BY-NC-SA 4.0 and our dataset Terms, the latter shall prevail. We once again emphasize that this dataset is exclusively intended for non-commercial purposes, such as academic research, teaching, or scientific publications. We strictly prohibit any commercial use of the dataset or any derived works, including the sale of data or utilization of data for commercial gain.